OpenStudy (anonymous):

How many solutions are there to the following system of equations? 2x + 6y = –12 10x + 32y = –62

OpenStudy (elonasushchik):

3. How many solutions are there to the following system of equations? 2x + 6y = –12 10x + 32y = –62 Answer: y = -x/3 - 2 y = -10x/32 - 62/32

OpenStudy (elonasushchik):

Medal and fan?! Thx

OpenStudy (anonymous):

thats not right

OpenStudy (mathstudent55):

Since the first equation has 2x and the second equation has 10x, multiply the first equation by 5 on both sides: 10x + 30y = -60 The second equation is still: 10x + 32y = -62 Since the two equations don't have the same x coefficients and y coefficients (just the same x coefficients), they have different slopes. Therefore, they must intersect at one single point, so there is one solution.